# Beyond Aesthetics: Cultural Competence in Text-to-Image Models

### v2

@@ 1 Introduction
In this paper, we present CUBE: CUltural BEnchmark, a first-of-its-kind benchmark designed to facilitate the evaluation of cultural competence of text-to-image (T2I) models along two axes: cultural awareness and cultural diversity. We build this benchmark at the country level (in line with other recent work (Jha et al., 2024; Li et al., 2024c)), encompassing eight countries and representing three different concepts of cultural artifacts chosen as concepts of clear visual elements and hence of importance to T2I models. For each of these concepts, we employed a large-scale extraction strategy that leverages a Knowledge Graph (KG) augmented with a large language model (LLM) to build a broad-coverage compilation of country-specific concepts/artifacts to ground our evaluation. CUBE consists of a) CUBE-1K - a carefully curated subset of 1000 artifacts, made into prompts that enable evaluation of cultural awareness (Nguyen et al., 2023) via human annotation, and b) CUBE-CSpace - a collection of  300K cultural artifacts spanning the 8 countries and 3 concepts we consider. The methodology used to collect this resource is scalable to other concepts and geo-cultures. Furthermore, we introduce cultural diversity (CD) as a new evaluation component for T2I models, adapting the quality-weighted Vendi score (Nguyen & Dieng, 2024).
@@ 3.2 CUBE-1K
As T2I models are primarily trained on English image-text pairs (Pouget et al., 2024), we expect them to struggle with visualizing long-tail artifacts from non-English-speaking cultures. To this end, CUBE-1K consists of prompts focusing on widely recognized artifacts, reflecting a model’s ability to capture mainstream cultural elements. The artifacts in CUBE-1K are a carefully curated subset of CUBE-CSpace. To ensure the inclusion of popular artifacts relevant to each country, we leverage the number of Google search results as a proxy for popularity. Specifically, we employ the Google Search API, utilizing the geolocation feature (’gl’ property) to tailor search results to a user located within the target country, thus capturing local popularity. We use this popularity estimate to sample artifacts for CUBE-1K. While search results serve as a useful proxy, we acknowledge they can be noisy, potentially inflated by the presence of popular keywords. Therefore, the final collection undergoes a manual verification process to ensure the quality and relevance of selected artifacts. CUBE-1K consists of 1000 prompts, spanning 8 countries and 3 concepts described above. Table 8 presents the distribution of artifacts across different countries in CUBE-1K. We then design prompt templates to probe the models for cultural awareness and specifically provide a negative prompt to obtain images of cultural artifacts with desirable qualities. Each prompt tests the T2I model’s ability to visualize a single cultural artifact. The prompt templates and negative prompt are provided in the Table 12.
@@ 4 Evaluating Cultural Awareness
To assess the cultural awareness of text-to-image (T2I) models, we leverage prompts from the CUBE-1K dataset. For this, we use traditional T2I evaluation aspects like faithfulness (adherence of the generated image to the input prompt) and realism (similarity of the generated image to a real photograph). These are measured using automated metrics like DSG (Cho et al., 2024) and FID (Heusel et al., 2018). However, these prove insufficient for capturing the complexities of cultural representation. Existing automated metrics are primarily trained on datasets lacking diverse cultural content and struggle to adequately assess the nuances of cultural elements. Therefore, we introduce a human annotation scheme specifically tailored to measure a model’s cultural awareness along two key dimensions: a) faithfulness and b) realism.
@@ 5.1.1 Foundation: Vendi Scores
Vendi scores are a family of interpretable diversity metrics that satisfy the axioms of ecological diversity (Friedman & Dieng, 2023; Pasarkar & Dieng, 2023). Vendi score captures the "effective number" of distinct items within a collection, considering both richness (number of unique elements) and evenness (distribution of those elements), and is defined as follows
@@
The order $q$ determines the sensitivity allocated to feature prevalence, with values of $q<1$ being more sensitive to rarer features and $q>1$ putting more emphasis on more common features. When $q=1$, we recover the original Vendi score (Friedman & Dieng, 2023), the exponential of the Shannon entropy of the normalized eigenvalues of $K$.
@@ 11 Acknowledgements
We thank Partha Talukdar, Kathy Meier-Hellstern, Caroline Pantofaru, Remi Denton, Susanna Ricco, David Madras, Nitish Gupta and Sunipa Dev for their feedback and advice; Lucas Beyer and Xiaohua Zhai for insights and support on use of mSigLIP for auto-evals; Sagar Gubbi and Kartikeya Badola for helpful discussions on the human rating template; Shikhar Vashishth for discussions on the use of WikiData; and Dinesh Tewari and the annotation team for facilitating our human evaluation work.

### v3

@@ 1 Introduction
In this paper, we present CUBE: CUltural BEnchmark, a first-of-its-kind benchmark designed to facilitate the evaluation of cultural competence of text-to-image (T2I) models along two axes: cultural awareness and cultural diversity. We build this benchmark at the country level (in line with other recent work (Jha et al., 2024; Li et al., 2024c)), encompassing eight countries and representing three different concepts of cultural artifacts chosen as concepts of clear visual elements and hence of importance to T2I models. For each of these concepts, we employed a large-scale extraction strategy that leverages a Knowledge Graph (KG) augmented with a large language model (LLM) to build a broad-coverage compilation of country-specific concepts/artifacts to ground our evaluation. CUBE consists of a) CUBE-1K - a carefully curated subset of 1000 artifacts, made into prompts that enable evaluation of cultural awareness (Nguyen et al., 2023) via human annotation, and b) CUBE-CSpace - a collection of  300K cultural artifacts spanning the 8 countries and 3 concepts we consider. The methodology used to collect this resource is scalable to other concepts and geo-cultures. Furthermore, we introduce cultural diversity (CD) as a new evaluation component for T2I models, adapting the quality-weighted Vendi score (nguyen2024qualityweighted).
@@ 3.2 CUBE-1K
As T2I models are primarily trained on English image-text pairs (Pouget et al., 2024), we expect them to struggle with visualizing long-tail artifacts from non-English-speaking cultures. To this end, CUBE-1K consists of prompts focusing on widely recognized artifacts, reflecting a model’s ability to capture mainstream cultural elements. The artifacts in CUBE-1K are a carefully curated subset of CUBE-CSpace. To ensure the inclusion of popular artifacts relevant to each country, we leverage the number of Google search results as a proxy for popularity. Specifically, we employ the Google Search API, utilizing the geolocation feature (’gl’ property) to tailor search results to a user located within the target country, thus capturing local popularity. We use this popularity estimate to sample artifacts for CUBE-1K. While search results serve as a useful proxy, we acknowledge they can be noisy, potentially inflated by the presence of popular keywords. Therefore, the final collection undergoes a manual verification process to ensure the quality and relevance of selected artifacts. CUBE-1K consists of 1000 prompts, spanning 8 countries and 3 concepts described above. Table 8 presents the distribution of artifacts across different countries in CUBE-1K. We use prompt templates designed to probe the models for cultural awareness, and specifically provide a negative prompt to obtain images of cultural artifacts with desirable qualities. Each prompt tests the T2I model’s ability to visualize a single cultural artifact. The prompt templates and negative prompt are provided in the Table 12.
@@ 4 Evaluating Cultural Awareness
To assess the cultural awareness of text-to-image (T2I) models, we leverage prompts from the CUBE-1K dataset. We use traditional T2I evaluation aspects like faithfulness (adherence of the generated image to the input prompt) and realism (similarity of the generated image to a real photograph) to measure cultural awareness. Conventionally, these are measured using automated metrics like DSG (Cho et al., 2024) and FID (Heusel et al., 2018). However, these prove insufficient for capturing the complexities of cultural representation. Existing automated metrics are primarily trained on datasets lacking diverse cultural content and struggle to adequately assess the nuances of cultural elements. Therefore, we introduce a human annotation scheme specifically tailored to measure a model’s cultural awareness along the two key dimensions: a) faithfulness and b) realism.
@@ 5.1.1 Foundation: Vendi Scores
Vendi scores are a family of interpretable diversity metrics that satisfy the axioms of ecological diversity (Dan Friedman & Dieng, 2023; Pasarkar & Dieng, 2023). Vendi score captures the "effective number" of distinct items within a collection, considering both richness (number of unique elements) and evenness (distribution of those elements), and is defined as follows
@@
The order $q$ determines the sensitivity allocated to feature prevalence, with values of $q<1$ being more sensitive to rarer features and $q>1$ putting more emphasis on more common features. When $q=1$, we recover the original Vendi score (Dan Friedman & Dieng, 2023), the exponential of the Shannon entropy of the normalized eigenvalues of $K$.
@@ 11 Acknowledgements
We thank Partha Talukdar, Kathy Meier-Hellstern, Caroline Pantofaru, Remi Denton, Susanna Ricco, David Madras, Nitish Gupta and Sunipa Dev for their feedback and advice; Lucas Beyer and Xiaohua Zhai for insights and support on use of mSigLIP for auto-evals; Sagar Gubbi and Kartikeya Badola for helpful discussions on the human rating template; Shikhar Vashishth for discussions on the use of WikiData; Preetika Verma for assistance with the SLING framework and Dinesh Tewari and the annotation team for facilitating our human evaluation work.